# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

### v1

@@ 3.2 Self-Improvement Dynamics
In this section, we aim to analyze the dynamics of self-improvement. Our techniques are inspired by the concept of potential energy, a widely used concept in physics (Rankine, 1853). Following Huang et al. (2025) and Song et al. (2025), we argue that the self-improvement comes from the Capability Gap $G(\cdot)$, defined as the gap between $U_{s}(t)$ and $U_{v}(t)$, namely,
@@
We assume that the change of the solver and verifier capability (Equation 6) is driven by a gap potential energy $E(t)$, assumed as a linear function of the Capability Gap $G(t)$ (Equation 8), namely
$E(t)=kG(t)-b,$ (9)
where $k,b$ denote the linear coefficient and bias. Experiments in Section 4 show that assuming potential energy in a linear form (as in Equation 9) is enough to explain empirical phenomena in self-improvement training. In this paper, we simply use uncertainty to evaluate the response quality, leading the capability gap defined with the uncertainty, as discussed in Section 3.1. We assume that the solver capability and the verifier capability both increase during the process of self-improvement, which is widely observed in the related works (Song et al., 2025). With the analysis above, following the concept of potential energy, the theoretical framework starts with the following assumptions:
@@
where $\alpha,\beta\geq 0$ are coefficients related to the decreasing rate of $U_{s}(t),U_{v}(t)$. In this framework, Equation 10 represents the initial conditions. Since for LLM, the verification capability usually outperforms the solver capability in real-world applications, we assume that $U_{s,0}>U_{v,0}$quation 11 represents the iterative conditions, where we assume that $\alpha>\beta$, indicating that the solver capability increases faster than the verifier. Based on the above assumptions, we derive the following dynamics:
$E(t)=k\gamma e^{-k(\alpha-\beta)t},$ $\quad G(t)=\gamma e^{-k(\alpha-\beta)t}+G_{\infty},$ (12)
$U_{s}(t)=\alpha^{\prime}e^{-k(\alpha-\beta)t}+U_{s,\infty},$ $\quad U_{v}(t)=\beta^{\prime}e^{-k(\alpha-\beta)t}+U_{v,\infty},$ (13)
where $\gamma=U_{s,0}-U_{v,0}-\frac{b}{k}$, $\alpha^{\prime}=\frac{\alpha}{\alpha-\beta}(U_{s,0}-U_{v,0}-\frac{b}{k})$, $\beta^{\prime}=\frac{\beta}{\alpha-\beta}(U_{s,0}-U_{v,0}-\frac{b}{k})$ represent coefficients, and $G_{\infty}=\frac{b}{k}$, $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$, $U_{v,\infty}=U_{v,0}-\beta^{\prime}$ represent the capability at convergence. We demonstrate from Equation 13 that $U_{s}$ and $U_{v}$ both decrease to an inherent uncertainty level obeying exponential law during self-improvement with $U_{s}$ decreasing much faster. The dynamics also indicate that (i) the gap potential energy $E(t)$ drives the solver capability stronger; (ii) the change of solver capability and verifier capability slows down as the gap decreases; and (iii) the capability gap might not necessarily converge to zero during the training process. We empirically validate our theoretical framework in Section 4.
@@ 4 Experiment
In this section, we conduct experiments on self-improvement to verify the theoretical framework. We first introduce experimental setups in Section 4.1. We then present the main experimental results in Section 4.2, sketching the dynamics of the self-improvement process. We finally explore the differences between the solver capability and the verifier capability, showing that the solver-verifier gap indeed generally happens in practice.
Figure 2: Accuracy and uncertainty during the self-improvement of the Phi4-mini model on the Math and GSM8k datasets using the QE method. The experimental results show that the accuracy increases during self-improvement process while the uncertainty decreases.
@@ 4.1 Setup
This section details the methods, models, datasets, and key parameters employed in our experiments. We consider the following verification methods: TrueFalse (TF): The solver generates $N$ responses, denoted $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier is then tasked with answering whether each response $\hat{y}_{i,j}$ is correct. If the verifier deems a response $\hat{y}_{i,j}$ correct, its score $s(\hat{y}_{i,j})$ is set to $1$; otherwise, it is set to 00; Quality Evaluation (QE): The solver generates $N$ responses, $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier then assigns a continuous score $s(\hat{y}_{i,j})$ between 0 and 1 to each response based on its quality. A score of $s(\hat{y}_{i,j})=0$ indicates a completely incorrect answer, while $s(\hat{y}_{i,j})=1$ indicates a completely correct answer.
We utilize two model families in our study: (a) Phi models: From the Phi family (Abdin et al., 2024), we use Phi4-Mini, Phi3.5-Mini, and Phi3-mini; (b) Llama models: We use Llama3.2-3B (Grattafiori et al., 2024) and Llama3.1-8B. Our experiments focus on the models’ mathematical problem-solving capabilities. Accordingly, we employ two representative datasets: GSM8k (Cobbe et al., 2021) and Math (Hendrycks et al., 2021). To maintain response diversity, we set the temperature to $1$, maximum length of $512$ tokens, threshold $\sigma=0.5$, and sample size $N=16$. Zero-shot prompting is used for all models and datasets. If the verifier cannot provide a definitive evaluation for an answer, we treat it as a neutral response and assign a score of $0.5$. Experimental details are provided in Appendix A.
@@ 4.2 Dynamics of Self-Improvement
This section sketches the dynamics of self-improvement using AI-generated feedback for supervised fine-tuning (SFT), aiming to verify the theoretical framework proposed in Section 3.2.
Setups. We tested the Phi4-mini, Phi3.5-mini, Phi3-mini, and Llama3.2-3B models. We applied Low-Rank Adaptation (LoRA) (Hu et al., 2022), which significantly reduces the number of updatable parameters, thereby enhancing SFT efficiency. The chosen hyperparameters are detailed in Section A.1. To improve training efficiency, we employed mini-batch gradient descent with a batch size of $256$. For each model-task pair, we conducted training for $10$ epochs, saving a checkpoint after processing half of the training data. We test the solver accuracy and the verifier accuracy, defined as the accuracy of the response and the BoN response, respectively.
@@
Results. We present the results of self-improvement on the Phi4-mini model using the QE method, as depicted in Figure 2. The empirical evidence indicates a consistent enhancement in the accuracy of both the solver and the verifier during the self-improvement process, coupled with a concurrent reduction in their respective uncertainties. Furthermore, a narrowing of the gap $G(t)$ between the solver and verifier is evident. Results for other models and methods are presented in Section A.2.
@@
Validation. To validate our theoretical framework and enable the prediction of the model’s final performance, we fit an exponential model to the uncertainty from $10$ self-improvement epochs. Figure 3 presents the results for the Phi4-mini model in four different settings (Math / GSM8K datasets with QE / TF metrics). The plots illustrate the evolution of three key metrics: solver uncertainty, verifier uncertainty, and the uncertainty gap. In all subfigures, the exponential model demonstrates a strong fit to the empirical data, with the coefficients of determination ($R^{2}$) exceeding $0.9$. This empirical evidence validates the exponential law proposed in our theoretical work.
Figure 3: Exponential trends of model uncertainty during self-improvement. The results illustrate the uncertainty associated with the Phi4-mini model’s solver and verifier, as well as their gap. The scatter points represent the measured data, while the solid lines are the best-fit curves to an exponential model. $R^{2}>0.9$ indicates that the evolution of these uncertainties is well-described by an exponential function.
Figure 4: Accuracy and uncertainty of Phi4-mini on Math and GSM8k with different sample size using TF and QE respectively. The results illustrate that the verifier perform better than the solver.
4.3 Verifier Outperform Solver
@@
In this section, we evaluate the solver and verifier performance of LLM, aiming to validate the utility of the solver-verifier gap used in Section 3. Specifically, the experimental results verify that verifier capability outperforms solver capability consistently. To evaluate the comprehensive capability of the model in the training set and test set, we randomly sample $2,048$ instances of the training set and $1,024$ instances from the test set for each data set. Our evaluation is divided into two settings: self-evaluation and cross-evaluation. Self-evaluation employs the same model for both the solver and verifier roles, whereas cross-evaluation utilizes different models for the solver and verifier.
@@
Self Evaluation. In this part, we compare the accuracy of solver and verifier on different models and datasets. Given that the verifier selects one response from $N$ candidates generated by the solver, the verifier’s performance is expected to improve as $N$ increases. For each model, we evaluate the accuracy and uncertainty of both the solver and the verifier, varying $N$ across the values ${2,4,8,16,32,64}$. As illustrated for the Phi4-mini model in Figure 4, an increase in $N$ corresponds to improved verifier accuracy and reduced uncertainty. This figure also shows that the verifier consistently outperforms the solver for Phi4-mini, achieving higher accuracy and lower uncertainty. Similar trends were observed across other tested models, with the possible exception of Phi4. Detailed results are presented in Section A.1.
Cross Evaluation. In self-evaluation, the same model serves as both solver and verifier. To better understand the relationship between solver and verifier capabilities, we also perform cross-evaluation, where one model acts as the solver and a different model acts as the verifier. These cross-evaluations utilize models with the QE method on both the MATH and GSM8k datasets. Furthermore, we set $N=16$ because, as demonstrated in Figure 4, accuracy does not improve significantly for $N>16$. The results of these evaluations are presented in Figure 5. We observe that when a fixed model serves as the solver, the verifier’s performance generally surpasses that of the solver, even when a different model is employed as the verifier.
@@
Figure 5: Cross-evaluation using QE method with sample size $N=16$ on Math. For each solver model, sampled responses are verified by different models. Figure on the left illustrates the difference of accuracy while figure on the right shows the $10$ logarithms of the uncertainty difference.
4.4 Pass@K
In this section, we investigate the underlying reason for the limit of self-improvement, focusing on how a decrease in the model’s response diversity leads to a diminishing potential energy, thereby causing the model’s capability to plateau. Although the self-improvement paradigm shows great improvement in model capability, it iteratively leads to a performance plateau. To investigate the underlying cause of this saturation, we conduct the Pass@K experiment on the initial model and the self-improved model. Pass@K is a metric that calculates the proportion of prompts where at least one of the $K$ responses is correct. When $K$ is large, Pass@K could be used to measure the diversity of the solver. We present the result of QE method in Figure 6. We observe that when $K$ is small, Pass@K increases with the number of epochs of self-improvement, validating the self-improvement process. However, when $K$ is large, we observe a slight decrease in Pass@K, indicating that the diversity of the solver is reduced through self-improvement. The degradation in diversity is caused by the convergence to a certain response, which is a potential reason for the limit of self-improvement.
Figure 6: Pass@K with QE method for different $K$ at $t=0$, $t=5$ and $t=10$. Pass@K evaluates the diversity of model generation. The efficacy of the self-improvement process is demonstrated by an increase in the Pass@K metric for small values of $K$. Conversely, for large values of $K$, a decrease in Pass@K is observed. This phenomenon suggests that the diversity of generations decreases during the self-improvement process.
@@ 5 Discussions on Cross-Improvement
In this section, we discuss how to model cross-improvement within the framework described above. The key insight is that external data affects the theoretical framework through the verification capability. We first present the theoretical framework in Section 5.1, under the framework with limited external data. We then conduct experiments in Section 5.2 to validate the theoretical findings.
@@ 5.1 Theoretical Framework of Cross-Improvement
In this section, we present the theoretical framework of training dynamics of cross-improvement. We follow the notations in Section 3. Besides, assume that we have limited external data with size $N^{\prime}$. For example, we may acquire $N^{\prime}$ external data in total from a better LLM using API queries.
We focus on the allocation of the external data. Specifically, for each epoch $t$, only $\eta_{t}N^{\prime}$ prompts could use API queries to get (one) external data, with $\sum_{t=1}^{T}\eta_{t}=1$ where $T$ denotes the total training epochs. For those chosen prompts, we choose the external data as the BoN response; for those non-chosen prompts, we still choose the BoN data from the $N$ internal responses. Notably, as long as one prompt is chosen to use external data, it will always use the external data.
@@
This assumption is a simplification of the effect of cross-improvement, and is demonstrated by experimental observations. The parameter $\gamma$ represents the general effect of cross-improvement, while $\eta_{t}$ represents the ratio of external data used in epoch $t$. We assume that the effect $\gamma$ is time-invariant, without which one cannot estimate the parameter. Overall, in each step, the cross-data first boosts the verification capability, then it evolves following the mathematical framework in Section 3.2. We illustrate this procedure in Figure 7.
@@
Equation 15 represents the initial conditions, which are the same as Equation 10Equation 16 represents the effects of cross-improvement, where the solver capability after cross-improvement $U_{s}^{c}(t)$ remains unchanged, while the verifier capability increases as discussed in Equation 14. The current capability gap $G^{c}(t)$ is then defined as the gap between the solver capability and the current verifier capability. Equation 18 represents the iterative conditions of cross-improvement. Note that Equation 18 differs slightly from the self-improvement dynamics in Equation 11 in the following ways: (i) we employ the external-data-affected state of solver capability $U_{s}^{c}(t)$ and verifier capability $U_{v}^{c}(t)$ during the process; (ii) we adopt discrete iteration instead of using a SDE. This is because formulating the SDE becomes challenging under the changes described in (i). Based on the above formulation, we derive an approximate solution of the ultimate uncertainty:
$\mathbf{U}(T)\approx e^{-\mathbf{\Delta^{\prime}}}\mathbf{U}(0),$ (19)
where $\mathbf{U}(T)$ denotes $\begin{bmatrix}U_{v,T},&U_{s,T},&1\end{bmatrix}^{\top}$ and $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. $U_{s}(T)$ is related only to the summation $\sum_{t=1}^{T}\eta_{t}$(instead of each $\eta_{t}$ at epoch $t$). Under the approximate solution, we come to the following conclusions:
  • For cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, the approximate solution is around the same;
  • The cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\sum_{t=1}^{T}\eta_{t}=0$ in terms of solver capability, when $\gamma>0$.
Detailed derivations of the above part are presented in Appendix B.

### v3

@@ 3.2 Self-Improvement Dynamics
In this section, we aim to analyze the dynamics of self-improvement. Given the immense complexity of the underlying neural network processes, deriving dynamics from first principles is currently intractable. Therefore, we adopt a phenomenological modeling approach, which is common in the study of complex systems. Our techniques are inspired by the concept of potential energy, a widely used concept in physics (Rankine, 1853). Following Huang et al. (2025) and Song et al. (2025), we argue that the self-improvement comes from the Capability Gap $G(\cdot)$, defined as the gap between $U_{s}(t)$ and $U_{v}(t)$, namely,
@@
We assume that the change of the solver and verifier capability is driven by a gap potential energy $E(t)$, expressed as a function of the Capability Gap $G(t)$ (Equation 7), namely $E(t)=f(G(t))$, where $f(G)$ is a differentiable and monotonically increasing function for $G\geq 0$, with its minimum $f(0)=0$. In this paper, we simply use uncertainty to evaluate the response quality, leading the capability gap defined with the uncertainty, as discussed in Section 3.1. We assume that the solver capability and the verifier capability both increase during the process of self-improvement, which is widely observed in the related works (Song et al., 2025). With the analysis above, following the concept of potential energy, the theoretical framework starts with the following assumptions:
@@
where $\alpha,\beta\geq 0$ are coefficients related to the decreasing rate of $U_{s}(t),U_{v}(t)$. In this framework, Equation 8 represents the initial conditions. Since for LLM, the verification capability usually outperforms the solver capability in real-world applications, we assume that $U_{s,0}>U_{v,0}$Equation 9 represents the iterative conditions, where we assume that $\alpha>\beta$, indicating that the solver capability increases faster than the verifier. Based on the above assumptions, we derive the differential equation governing the Capability Gap:
$\frac{dG(t)}{dt}=-(\alpha-\beta)E(t).$ (10)
However, for a non-linear function $f(G)$, a closed-form solution for above integral is not guaranteed. To simplify this problem, we approximate the potential energy function $E(t)$ with a linear form, $E(t)\approx kG(t)-b$, derived from a first-order Taylor expansion. Figure 2 illustrates the strong linear relationship between uncertainty gap $G$ and its rate of change $dG/dt=-(\alpha-\beta)E(t)$ on Phi-4-mini with QE method, which provides robust evidence for our approximation. Experiment results on other models are presented in Figure 7. Based on the Equation 9, we derive proposition 3.1:
Figure 2: Verification on linear relationship between the Uncertainty gap $G$ and its rate of change $dG/dt$ on Phi-4-mini with QE method. The scatter points represent empirical data from self-improvement on the Math and GSM8k datasets, while the solid lines show the linear regression fits.
Proposition 3.1.
Assume $k(\alpha-\beta)>0$, the dynamics of the potential energy $E(t)$, the capability gap $G(t)$. Let the potential energy function $f(G)$ be linearly approximated around an initial state $G_{0}$ by its tangent line $f(G)\approx kG-b$, where $k=f^{\prime}(G_{0})$ and $b=f^{\prime}(G_{0})G_{0}-f(G_{0})$. Capabilities $U_{s}(t)$ and $U_{v}(t)$ are governed by the following exponential decay functions:
$E(t)\approx k\delta e^{-k(\alpha-\beta)t},$ $\quad G(t)\approx\delta e^{-k(\alpha-\beta)t}+G_{\infty},$ (11)
$U_{s}(t)\approx\alpha^{\prime}e^{-k(\alpha-\beta)t}+U_{s,\infty},$ $\quad U_{v}(t)\approx\beta^{\prime}e^{-k(\alpha-\beta)t}+U_{v,\infty},$ (12)
where the coefficients are defined as: $\delta=U_{s,0}-U_{v,0}-\frac{b}{k},\alpha^{\prime}=\frac{\alpha\delta}{\alpha-\beta},\beta^{\prime}=\frac{\beta\delta}{\alpha-\beta}$, and the values at convergence ($t\to\infty$) are given by: $G_{\infty}=\frac{b}{k},U_{s,\infty}=U_{s,0}-\alpha^{\prime},U_{v,\infty}=U_{v,0}-\beta^{\prime}$.
Equation 11 and Equation 12 demonstrate that uncertainties $U_{s}$, $U_{v}$, and their gap $G(t)$ all converge exponentially toward their respective limits. The proof of proposition 3.1 is shown in Section B.1. We derive two key corollaries from proposition 3.1.
Corollary 3.1.
The partial derivative of the solver’s final uncertainty with respect to the initial gap is a negative constant: $\frac{\partial U_{s,\infty}}{\partial G_{0}}=-\frac{\beta}{\alpha-\beta}.$
The proof of corollary 3.1 is shown in Section B.2corollary 3.1 shows that a larger initial verifier-solver gap ($G_{0}$) leads to a lower final solver uncertainty ($U_{s,\infty}$). This aligns with the core intuition that the performance gap is the driver of the self-improvement process.
Corollary 3.2.
For a given tolerance $\epsilon>0$, the training time $t$ required to ensure the Capability Gap is within $\epsilon$ of its convergence limit satisfies: $t>\frac{\ln(\delta/\epsilon)}{k(\alpha-\beta)}.$
The proof of corollary 3.2 is shown in Section B.3corollary 3.2 provides theoretical guidance for selecting the total number of training epochs $T$ in practical application. In practice, we first estimate the key convergence parameters by training for a few initial epochs and fitting the early-stage performance data. These parameters can then be used to determine an appropriate total training time $T$ to balance performance with computational costs.
The dynamics indicate that (i) the gap potential energy $E(t)$ drives the solver capability more strongly; (ii) the change of solver capability and verifier capability slows down as the gap decreases; and (iii) the capability gap might not converge to zero during the training process.
@@ 4 Experiment
This section conducts experiments on self-improvement to verify the theoretical framework. We first introduce experimental setups in Section 4.1. We then present experiment results of the self-improvement process in Section 4.2. We finally explore the differences between solver capability and verifier capability in Section 4.3, showing that the solver-verifier gap generally happens in practice.
Figure 3: Accuracy and uncertainty during the self-improvement of the Phi-4-mini model on the Math and GSM8k datasets using the QE method. The experimental results show that the accuracy increases during self-improvement process while the uncertainty decreases.
@@ 4.1 Setup
This section introduces the models, datasets, and key parameters. We consider the following two verification methods: TrueFalse (TF) and Quality Evaluation (QE). We utilize two model families: (a) Phi models: From the Phi family (Abdin et al., 2024), we use Phi-4-mini, Phi-3.5-mini, and Phi-3-mini; (b) Llama models: We use Llama-3.2-3B (Grattafiori et al., 2024) and Llama-3.1-8B. Our experiments mainly focus on the models’ mathematical problem-solving capabilities. Accordingly, we employ two representative datasets: GSM8k (Cobbe et al., 2021) and Math (Hendrycks et al., 2021). In addition, we also consider ProntoQA dataset with TF method in order to evaluate the efficiency of our framework on QA problems. Experimental details are provided in Appendix C.
@@ 4.2 Dynamics of Self-Improvement
This section sketches the dynamics of self-improvement using AI-generated feedback for supervised fine-tuning (SFT), aiming to verify the theoretical framework proposed in Section 3.2. We test the Phi-4-mini, Phi-3.5-mini, Phi-3-mini, and Llama-3.2-3B models. We apply Low-Rank Adaptation (LoRA) (Hu et al., 2022), which significantly reduces the number of updatable parameters, thereby enhancing SFT efficiency. The chosen hyperparameters are detailed in Section C.1.
@@
Results. We present the results of self-improvement on the Phi-4-mini model using the QE method, as depicted in Figure 3. The empirical evidence indicates a consistent enhancement in the accuracy of both the solver and the verifier during the self-improvement process, coupled with a concurrent reduction in their respective uncertainties. Furthermore, a narrowing of the gap $G(t)$ between the solver and verifier is evident. Results for other models and methods are presented in Section C.2.
@@
Validation. To validate our theoretical framework, we fit an exponential model to the uncertainty from $10$ self-improvement epochs. Figure 4 presents the results for the Phi-4-mini model in four different settings (Math / GSM8K datasets train split with QE / TF metrics), which illustrates the evolution of three key metrics: solver uncertainty, verifier uncertainty, and the uncertainty gap. In Figure 4, the exponential model demonstrates a strong fit to the empirical data, with the coefficients of determination ($R^{2}$) exceeding $0.9$. This empirical evidence validates the exponential law proposed in our theoretical work. Results for other models are presented in Section C.2.
4.3 Verifier Outperforms Solver
@@
In this section, we evaluate the solver and verifier performance of LLM, aiming to validate the utility of the solver-verifier gap used in Section 3. Specifically, the experimental results verify that verifier capability outperforms solver capability consistently. Our evaluation is divided into two settings: self-evaluation and cross-evaluation. Self-evaluation employs the same model for both the solver and verifier roles, whereas cross-evaluation utilizes different models for the solver and verifier.
@@
Self Evaluation. We compare the accuracy and uncertainty of solver and verifier on different models and datasets across different $N$. Given that the verifier selects one response from $N$ candidates generated by the solver, the verifier’s performance is expected to improve as $N$ increases. Detailed results are presented in Section C.2.
Cross Evaluation. To better understand the relationship between solver and verifier capabilities, we also perform cross-evaluation, where one model acts as the solver and a different model acts as the verifier. Results are presented in Section C.2.
Figure 4: Exponential trends of model uncertainty during self-improvement on train split. The results illustrate the uncertainty and gap of the solver and verifier for Phi-4-mini. The scatter points represent the measured data, while the solid lines are the best-fit curves to an exponential model. $R^{2}>0.9$ indicates that the evolution of these uncertainties is well-described by an exponential function.
@@ 5 Discussions on Cross-Improvement
In this section, we discuss how to model cross-improvement within the framework described in Section 3.2. The key insight is that external data affects the theoretical framework through the verification capability. We first present the theoretical framework in Section 5.1, under the framework with limited external data. We then conduct experiments in Section 5.2 to validate the theoretical findings.
@@ 5.1 Theoretical Framework of Cross-Improvement
In this section, we present the theoretical framework of training dynamics of cross-improvement. We follow the notations in Section 3. Besides, assume that we have limited external data with size $M$. For example, we may acquire $M$ external data in total from a better LLM using API queries.
We focus on the allocation of the external data. Specifically, for each epoch $t$, only $\eta_{t}M$ prompts could use API queries to get (one) external data, with $\sum_{t=1}^{T}\eta_{t}=1$ where $T$ denotes the total training epochs. For those chosen prompts, we choose the external data as the BoN response; for those non-chosen prompts, we still choose the BoN data from the $N$ internal responses. Notably, as long as one prompt is chosen to use external data, it will always use the external data.
@@
The parameter $\gamma$ represents the general effect of cross-improvement, while $\eta_{t}$ represents the ratio of external data used in epoch $t$. We assume that the effect $\gamma$ is time invariant, $\gamma>0$ when the verifier that adds external data performs better than the verifier without external data; otherwise $\gamma<0$. Overall, in each step, the cross-data first boosts the verification capability, then it evolves following the mathematical framework in Section 3.2. We illustrate this procedure in Figure 6.
@@
Equation 14 represents the initial conditions, which are the same as Equation 8Equation 15 represents the effects of cross-improvement, where the solver capability after cross-improvement $U_{s}^{c}(t)$ remains unchanged, while the verifier capability increases as discussed in Equation 13. The current capability gap $G^{c}(t)$ is then defined as the gap between the solver capability and the current verifier capability. Equation 17 represents the iterative conditions of cross-improvement. Based on the above formulation, we derive an approximate solution of the ultimate uncertainty:
Proposition 5.1.
Let the external-data-affected solver uncertainty $U_{s}^{c}(t)$ and verifier uncertainty $U_{v}^{c}(t)$. The ultimate uncertainty vector $\mathbf{U}(T)$ can be approximated as
$\mathbf{U}(T)\approx e^{-\mathbf{\Delta^{\prime}}}\mathbf{U}(0),\quad\mathbf{\Delta^{\prime}}=\begin{pmatrix}T-(1+\beta k)(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\beta k&-T\beta b\\ -\alpha k(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\alpha k&-T\alpha b\\ 0&0&0\end{pmatrix}$, (18)
where $\mathbf{U}(T)$ denotes $\begin{bmatrix}U_{v,T},&U_{s,T},&1\end{bmatrix}^{\top}$ and $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. $U_{s}(T)$ is related only to the summation $\sum_{t=1}^{T}\eta_{t}$ (instead of each $\eta_{t}$ at epoch $t$).
Under proposition 5.1, we come to the following conclusions: (i) For cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, the approximate solution is around the same; (ii) The cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\sum_{t=1}^{T}\eta_{t}=0$ in terms of solver capability, when $\gamma>0$. The proof of proposition 5.1 is shown in Section B.4.